### v1

# Gluon Distributions in the Pion

Abstract
We formulate a light-front model for the pion that explicitly incorporates the gluonic degree of freedom. In this framework, high-energy scattering off the pion is described by an active gluon, while the remaining constituents are treated as a spectator system with an effective mass. The mass spectrum and light-front wave functions (LFWFs) of the pion are determined by solving two Schrödinger-like equations derived from quantum chromodynamics: the light-front holographic equation in the chiral limit and the ’t Hooft equation. Using the resulting LFWFs, we compute the gluon distributions within pion, including the parton distribution functions in comparison with the available global fits, generalized parton distributions, and transverse momentum-dependent distributions. Furthermore, we demonstrate that the obtained LFWFs provide a good description of the gravitational form factors of pion when compared with the recent lattice QCD results.
I Introduction
Quantum Chromodynamics (QCD) is the fundamental theory of the strong force Callan et al. (1978), describing hadrons as bound states of quarks and gluons. Despite its successes, key aspects of hadron structure—particularly the role of gluons—remain poorly understood. Gluons are essential for binding quarks and generating mass, yet their distribution inside light mesons such as the pion is still not well-determined.
The pion plays a unique role in QCD: it is the lightest meson, a quark-antiquark bound state, and the Nambu-Goldstone boson of spontaneously broken chiral symmetry Nambu and Jona-Lasinio (1961a, b). Its simplicity makes it an ideal system for probing nonperturbative QCD. While most theoretical studies have focused on valence quark structure or incorporated gluons only perturbatively de Melo et al. (2009); Frederico et al. (2009); Chang et al. (2013); Pasquini and Schweitzer (2014); Gutsche et al. (2015); Lorcé et al. (2016); Bacchetta et al. (2017); de Teramond et al. (2018); Ahmady et al. (2018, 2020, 2019a); Raya et al. (2022); Lan et al. (2019); de Paula et al. (2022); Lu et al. (2022); Cui et al. (2022); Mondal et al. (2021); Adhikari et al. (2021); Ahmady et al. (2021a, b); Lan and Mondal (2020), recent advances in lattice QCD Fan and Lin (2021); Shanahan and Detmold (2019) and phenomenological models Pasquini et al. (2023); Lan et al. (2022); Zhu et al. (2023); Lan et al. (2025); Broniowski et al. (2023) now explicitly account for gluon dynamics, underscoring their importance in pion structure.
Light-front quantization offers a powerful framework for studying hadrons through light-front wave functions (LFWFs) Brodsky et al. (1998), which encode partonic configurations and enable the calculation of hadronic matrix elements for both inclusive and exclusive processes Lorce et al. (2011). This approach is well suited for investigating the partonic structure of hadrons.
In this work, we compute gluon-related observables in the pion including the parton distribution functions (PDFs), generalized parton distributions (GPDs), and transverse momentum dependent distributions (TMDs). PDFs describe longitudinal momentum distributed among partons, while GPDs extend this to spatial and momentum correlations, depending on the momentum fraction $x$, skewness $\xi$, and momentum transfer $t=\Delta^{2}$. Their moments relate to gravitational form factors (GFFs), revealing insights into hadron mass, pressure, and energy distributions Ji (1997); Polyakov and Schweitzer (2018); Lorcé et al. (2019); Burkert et al. (2023). TMDs further enrich this picture by resolving transverse parton momenta.
Global analyses of pion PDFs, primarily using Drell-Yan, $J/\psi$ production, and direct photon data Barry et al. (2021); Novikov et al. (2020); Pasquini et al. (2023), have advanced our understanding of its structure. However, the gluon and sea-quark content remains less constrained than in the nucleon due to limited experimental data. Upcoming facilities like Jefferson Lab Arrington et al. (2022) and future electron-ion colliders (EICs) Abdul Khalek et al. (2022); Anderle et al. (2021) aim to probe pion PDFs via the Sullivan process Sullivan (1972). The COMPASS++/AMBER program will further explore pion structure using high-intensity pion beams Adams et al. (2018).
Experimental access to GPDs is challenging but progressing through deeply virtual Compton scattering (DVCS) and vector meson production Diehl (2023); Adloff et al. (2000); Airapetian et al. (2001), with data from H1, ZEUS, and HERMES Adloff et al. (2001); Aktas et al. (2005); Aaron et al. (2009); Chekanov et al. (2003); Airapetian et al. (2010, 2012a, 2012b) already providing constraints. Near-threshold photo- and leptoproduction of charmonium and bottomonium provides a powerful probe of the hadron’s gluonic structure Mamo and Zahed (2020, 2022); Guo et al. (2024); Hatta and Yang (2018); Boussarie and Hatta (2020); Pentchev and Chudakov (2024); Ali et al. (2019); Adhikari et al. (2023). Future EICs are expected to extend these studies Abdul Khalek et al. (2022); Anderle et al. (2021), while Jefferson Lab has already reported promising results from near-threshold $J/\psi$ production Ali et al. (2019); Duran et al. (2023); Adhikari et al. (2023); Liu and Zahed (2024); Hechenberger et al. (2024).
While TMDs have been extensively studied for the proton via SIDIS, DY, and $e^{+}e^{-}$ annihilation Diehl (2023), pion TMDs remain less explored, with only limited extractions from DY data Wang et al. (2017); Vladimirov (2019); Cerutti et al. (2023). Theoretical investigations of pion TMDs have been pursued in Refs. Pasquini and Schweitzer (2014); Noguera and Scopetta (2015); Lorcé et al. (2016); Bacchetta et al. (2017); Ceccopieri et al. (2018); Ahmady et al. (2019a); Shi et al. (2020); Kaur et al. (2020), but further experimental and theoretical efforts are required.
In this work, we develop a light-front model of the pion that explicitly incorporates gluonic degrees of freedom, providing a framework to explore the pion’s gluonic structure. In this approach, high-energy scattering processes involving the pion are modeled as being mediated by an active gluon, while the remaining constituents form an effective spectator system characterized by a dynamically generated mass. The mass and LFWFs of the pion are determined self-consistently by solving two complementary equations rooted in QCD: the light-front holographic Schrödinger equation, which captures the transverse dynamics and confinement mechanism in the chiral limit Brodsky et al. (2015), and the ’t Hooft equation, which governs the longitudinal momentum structure of mesons in the large-$N_{c}$ limit ’t Hooft (1974). Together, these equations offer a three-dimensional description of the pion’s bound-state dynamics, incorporating both the transverse and longitudinal aspects of confinement. This unified framework provides predictions for gluon-sensitive observables, including the gluon PDFs, GPDs, TMDs, and GFFs of the pion. The model predictions show good agreement with recent lattice QCD results and the available global fits, demonstrating its effectiveness in capturing both the momentum and spatial structure of the pion.
II Theoretical framework
Holographic light-front QCD (hLFQCD) is formulated in the chiral limit of light-front QCD and establishes a correspondence between strongly coupled (3+1)-dimensional light-front QCD and weakly coupled string modes in a (4+1)-dimensional anti-de Sitter (AdS) space. For a comprehensive review of hLFQCD, see Ref. Brodsky et al. (2015). A central prediction of hLFQCD is that the pion, as the lightest bound state, becomes massless in the chiral limit. Moreover, the model predicts that meson masses lie on universal Regge trajectories, in agreement with experimental observations. The slopes of these trajectories are governed by the strength of the confining potential, characterized by the parameter $\kappa$. In this framework, the confining potential in physical spacetime emerges from a dilaton field that breaks the conformal symmetry of AdS space. Specifically, a quadratic dilaton profile in the fifth dimension leads to a light-front harmonic oscillator potential, which has proven to be phenomenologically successful. The scale parameter $\kappa$ is fixed by fitting the experimentally measured slopes of Regge trajectories across various meson families, yielding a universal value of $\kappa\simeq 0.5$ GeV for light mesons Brodsky et al. (2015).
The minimal Fock state of a meson consists exclusively of a valence quark-antiquark pair. When higher Fock sectors are included, the meson wavefunction acquires additional components involving gluons and sea quark-antiquark pairs. In our simplified model, we describe the meson as a composite system made up of an active gluon and an effective spectator. The LFWF $\Psi(x,\vec{b}_{\perp})$ provides a Lorentz-invariant description of the gluon-spectator dynamics in terms of the gluon’s longitudinal momentum fraction, $x=k^{+}/P^{+}$, where $k^{+}$ and $P^{+}$ denote the light-front momenta of the gluon and the meson, respectively. The variable $\vec{b}_{\perp}=b_{\perp}e^{i\varphi}$ represents the transverse separation between the gluon and the spectator. To establish a connection with AdS space, the holographic variable
$\vec{\zeta}=\sqrt{x(1-x)}\,\vec{b}_{\perp}=\zeta e^{i\varphi}$ (1)
is introduced. This allows the LFWF to be expressed in a factorized form in terms of $x$, $\zeta$, and $\varphi$ as
$\Psi(x,\zeta,\varphi)=\frac{\phi(\zeta)}{\sqrt{2\pi\zeta}}\,e^{iL\varphi}\,X(x)\,,$ (2)
where $\phi(\zeta)$ and $X(x)=\sqrt{x(1-x)}\,\chi(x)$ represent the transverse and longitudinal modes, respectively.
The hLFQCD, defined in the chiral limit of QCD and containing only the dynamical transverse mode, is governed by a holographic Schrödinger-like equation Brodsky and de Téramond (2006); de Téramond and Brodsky (2005, 2009); Brodsky et al. (2015):
$\left(-\frac{\mathrm{d}^{2}}{\mathrm{d}\zeta^{2}}+\frac{4L^{2}-1}{4\zeta^{2}}+U^{\mathrm{hLF}}_{\perp}(\zeta)\right)\phi(\zeta)=M_{\perp}^{2}\phi(\zeta)\;,$ (3)
where the transverse confinement potential is given by
$U_{\perp}^{\mathrm{hLF}}(\zeta)=\kappa^{4}\zeta^{2}+2\kappa^{2}(J-1)\;,$ (4)
with $J=L+S$ (where $L\equiv|L_{z}^{\mathrm{max}}|$) denoting the total angular momentum of the meson. The functional form of $U_{\perp}^{\mathrm{hLF}}(\zeta)$ is uniquely determined Brodsky et al. (2014) by the underlying conformal symmetry and the holographic correspondence to $\mathrm{AdS}_{5}$, where the light-front variable $\zeta$ is identified with the fifth dimension of AdS space. The mass scale parameter $\kappa$ defines both the confinement strength and the meson mass spectrum in the chiral limit. The Schrödinger equation (3) can be solved analytically, yielding the mass eigenvalues
$M_{\perp}^{2}(n_{\perp},J,L)=4\kappa^{2}\left(n_{\perp}+\frac{J+L}{2}\right)\;,$ (5)
and the corresponding transverse wavefunctions
$\phi_{n_{\perp}L}(\zeta)\propto\zeta^{1/2+L}\exp\left(-\frac{\kappa^{2}\zeta^{2}}{2}\right)L_{n_{\perp}}^{L}(\kappa^{2}\zeta^{2})\;.$ (6)
A notable prediction of Eq. (5) is that the lowest-lying bound state, corresponding to $n_{\perp}=L=S=0$, is massless. This naturally identifies the pion as the chiral-limit Goldstone boson in QCD.
In contrast, the longitudinal mode $\chi(x)$ is not dynamical in hLFQCD and is typically set to a constant, $\chi(x)=1$, leading to $X(x)=\sqrt{x(1-x)}$. This choice is validated by the holographic mapping of the pion’s electromagnetic and gravitational form factors between physical spacetime and $\mathrm{AdS}_{5}$ Brodsky and de Teramond (2008a, b).
Substituting Eq. (6) into Eq. (2) yields the chiral-limit LFWF for the pion:
$\Psi^{\pi}_{c}(x,\zeta^{2})\propto\sqrt{x(1-x)}\exp\left(-\frac{\kappa^{2}\zeta^{2}}{2}\right)\;.$ (7)
Applying a two-dimensional Fourier transform to Eq. (7) gives the corresponding momentum-space LFWF:
$\Phi^{\pi}_{c}(x,k^{2}_{\perp})\propto\frac{1}{\sqrt{x(1-x)}}\exp\left(-\frac{{k}_{\perp}^{2}}{2\kappa^{2}x(1-x)}\right)\;,$ (8)
where $\vec{k}_{\perp}$ denotes the transverse momentum of the quark, Fourier-conjugate to the transverse separation ${\vec{b}_{\perp}}$.
Beyond the chiral limit, Brodsky and de Téramond proposed a prescription for the longitudinal mode based on the invariant mass Brodsky and de Téramond (2009). While this approach has two notable limitations—it does not satisfy the Gell-Mann-Oakes-Renner (GMOR) relation and assigns the same longitudinal mode to all radially excited states—it has nonetheless been successfully applied to describe various properties of both light and heavy mesons Gurjar et al. (2024); Brodsky et al. (2015); Brodsky and de Téramond (2009); de Téramond and Brodsky (2021); Swarnkar and Chakrabarti (2015); Ahmady et al. (2020, 2019b, 2019a, 2018, 2017). An alternative approach to modeling the longitudinal dynamics uses the ’t Hooft equation ’t Hooft (1974) from large-$N_{c}$ QCD in (1+1) dimensions, which has been shown to reproduce the full meson mass spectrum, including the pion Ahmady et al. (2023), $\rho$ meson Gurjar et al. (2024), and both heavy-light and heavy-heavy systems Ahmady et al. (2021b). The idea of replacing the invariant mass prescription with the ’t Hooft equation was first introduced in Ref. Chabysheva and Hiller (2013), which focused on predicting meson decay constants and parton distribution functions. In parallel, recent studies de Téramond and Brodsky (2021); Li and Vary (2022) have employed a phenomenological longitudinal confinement potential, initially proposed in Ref. Li et al. (2016) within the basis light-front quantization (BLFQ) framework. Although these works primarily address the chiral limit and chiral symmetry breaking, Ref. de Téramond and Brodsky (2021) also explores applications to ground-state heavy mesons and connects their approach to the ’t Hooft equation. These developments reflect a growing interest in incorporating longitudinal dynamics into hLFQCD Li and Vary (2022); de Téramond and Brodsky (2021); Lyubovitskij and Schmidt (2022); Weller and Miller (2022); Ahmady et al. (2021a); Rinaldi et al. (2022).
Starting from the QCD Lagrangian in $(1+1)$ dimensions and applying the large-$N_{c}$ approximation, ’t Hooft derived a Schrödinger-like equation for the meson wavefunction ’t Hooft (1974):
$\left(\frac{m_{1}^{2}}{x}+\frac{m_{2}^{2}}{1-x}\right)\chi(x)+U_{\parallel}(x)\chi(x)=M^{2}_{\parallel}\chi(x)\;,$ (9)
where $m_{1}$ and $m_{2}$ are the masses of constituents and the longitudinal potential is given by
$U_{\parallel}(x)\chi(x)=\frac{g^{2}}{\pi}\,\mathcal{P}\int{\rm d}y\,\frac{\chi(x)-\chi(y)}{(x-y)^{2}}\;,$ (10)
with $g$ denoting the longitudinal confinement scale and $\mathcal{P}$ representing the Cauchy principal value. Unlike the holographic light-front Schrödinger equation, the ’t Hooft equation does not admit an analytical solution and must be solved numerically. Here, we follow the matrix method outlined in Ref. Chabysheva and Hiller (2013).
Combining the holographic Schrödinger equation for the transverse mode with the ’t Hooft equation for the longitudinal mode, the meson mass spectrum is given by:
$M^{2}(n_{\perp},n_{\parallel},J,L)=4\kappa^{2}\left(n_{\perp}+\frac{J+L}{2}\right)+M^{2}_{\parallel}(n_{\parallel},m_{1},m_{2},g)\;,$ (11)
where $n_{\parallel}$ is the longitudinal quantum number. For the ground state ($n_{\perp}=L=S=n_{\parallel}=0$), this formula reproduces the physical pion mass, in contrast to Eq. (5), where the pion is massless in the chiral limit. This shows that the ’t Hooft equation accounts for the full mass of the pion.
Table 1: Quantum numbers and masses of the pion family.
$J^{P(C)}$ Name $n_{\perp}$ $n_{\parallel}$ $L$ $M_{\perp}$ (MeV) $M_{\parallel}$ (MeV) $M$ (MeV)
$0^{-}$ $\pi(140)$ 0 0 0 0 135 135
$0^{-+}$ $\pi(135)$ 0 0 0 0 135 135
$1^{+-}$ $b_{1}(1235)$ 0 2 1 1046 303 1089
$0^{-+}$ $\pi(1300)$ 1 2 0 1046 303 1089
$2^{-+}$ $\pi_{2}(1670)$ 0 4 2 1479 409 1534
$0^{-+}$ $\pi(1800)$ 2 4 0 1479 409 1534
$2^{-+}$ $\pi_{2}(1880)$ 1 6 2 1811 492 1877
III Pion Mass Spectroscopy and Light-Front Wavefunctions
We employ a physical gluon mass, setting $m_{1}=0$. To fit the spectroscopic data of the pion family, we fix the three free parameters of our model: the longitudinal confinement scale $g$, the transverse confinement scale $\kappa$, and the spectator mass $m_{2}$. We take $m_{2}=0.092$ GeV, which is twice the light quark mass used in hLFQCD with the Brodsky–de Téramond ansatz Brodsky et al. (2015). For the transverse confinement scale, we use the universal value $\kappa=0.523$ GeV, consistent across the full hadron spectrum Ahmady et al. (2021b). In contrast, the longitudinal confinement scale $g$ is not universal; here, we use $g=0.109$ GeV Ahmady et al. (2021b). Following Refs. Ahmady et al. (2021a, b, 2023), the parity and charge conjugation quantum numbers of mesons are given by
$P=(-1)^{L+1},\quad C=(-1)^{L+S+n_{\parallel}}.$ (12)
We present the computed masses of the pion and its excitations in Table 1. Our results (last column) show good agreement with experimental values (second column, in parentheses). Notably, the condition $n_{\parallel}\geq n_{\perp}+L$ Ahmady et al. (2021b), observed in Table 1, appears to hold throughout the hadron spectrum. The corresponding Regge trajectories for the pion family are shown in Fig. 1.
Figure 1: Our Regge trajectories for the pion family using the physical gluon mass $m_{1}=0$, the spectator mass $m_{2}=0.092$ GeV, $g=0.109$ GeV and $\kappa=0.523$ GeV.
Although the longitudinal wavefunction $\chi(x)$ cannot be determined analytically, it can be well approximated by:
$\chi(x)\approx a_{0}+a_{1}x+a_{2}x^{2}+a_{3}x^{3}+a_{4}x^{4}+a_{5}x^{5}\;,$ (13)
where $a_{i}$ are constituent-mass-dependent parameters that vanish in the chiral limit except $a_{0}(=1)$. With the chosen parameters, we find that for the pion ground state, $\{a_{0},a_{1},a_{2},a_{3},a_{4},a_{5}\}\approx\{1.57,-0.49,-2.47,1.53,0.50,-0.62\}$. Figure 2 illustrates the dynamical longitudinal wavefunction of the ground-state pion: $X(x)=\sqrt{x(1-x)}\chi(x)$, where $x$ is momentum fraction carried by the gluon.
Incorporating both the transverse mode from the light-front Schrödinger equation and the longitudinal mode from the ’t Hooft equation, the spin-independent LFWF of the meson takes the form:
$\Phi^{\pi}(x,{k}^{2}_{\perp})=\mathcal{N}\frac{\chi(x)}{\sqrt{x(1-x)}}\exp\left(-\frac{{k}_{\perp}^{2}}{2\kappa^{2}x(1-x)}\right)\;,$ (14)
where $\chi(x)$ is the longitudinal mode obtained from the ’t Hooft equation and $\mathcal{N}$ is a normalization constant.
Figure 2: The longitudinal wave function $X(x)$ of gluon in the pion.
In our gluon-spectator model, we assume that the spectator has spin 1, same as the gluon. The two-particle Fock-state expansion for the pion’s orbital angular momentum component, $L^{z}=0$, in a frame where the pion’s transverse momentum vanishes, i.e., $P\equiv\big{(}P^{+},\vec{0}_{\perp},\frac{M^{2}}{P^{+}}\big{)}$, is expressed as Pasquini et al. (2023)
$|\pi(P)\rangle=$ $\int\frac{\mathrm{d}^{2}\vec{k}_{\perp}\mathrm{d}x}{16\pi^{3}\sqrt{x(1-x)}}\Big{[}\psi_{-1,+1}^{\pi}\left(x,\vec{k}_{\perp}\right)\left|-1,+1;xP^{+},\vec{k}_{\perp}\right\rangle$
$+\psi_{+1,-1}^{\pi}\left(x,\vec{k}_{\perp}\right)\left|+1,-1;xP^{+},\vec{k}_{\perp}\right\rangle\Big{]}.$ (15)
For nonzero transverse momentum of the pion, i.e., $\vec{P}_{\perp}\neq 0$, the physical transverse momenta of the gluon and spectator are given by $\vec{p}_{\perp}^{\,g}=x\vec{P}_{\perp}+\vec{k}_{\perp}$ and $\vec{p}_{\perp}^{\,s}=(1-x)\vec{P}_{\perp}-\vec{k}_{\perp}$, respectively, where $\vec{k}_{\perp}$ represents the relative transverse momentum of the gluon. The LFWFs, denoted by $\psi_{\lambda_{g},\lambda_{s}}^{\pi}(x,\vec{k}_{\perp})=\Phi^{\pi}(x,\vec{k}_{\perp})\lambda_{g}\delta_{\lambda_{g},-\lambda_{s}}$ correspond to the two-particle state $|\lambda_{g},\lambda_{s};xP^{+},\vec{k}_{\perp}\rangle$, with the gluon’s helicity $\lambda_{g}=\pm 1$, and the spectator’s helicity $\lambda_{s}=\pm 1$. The LFWF is normalized as follows:
$\sum_{\lambda_{g},\lambda_{s}}\int\frac{{\mathrm{d}}x\,{\mathrm{d}}^{2}\vec{k}_{\perp}}{16\pi^{3}}\left|\psi^{\pi}_{\lambda_{g},\lambda_{s}}(x,\vec{k}_{\perp})\right|^{2}=1\;.$ (16)
IV Gluon generalized parton distributions
The pion gluon GPDs are defined through the off-forward matrix elements of the bilocal gluon tensor operator between pion states Diehl (2003):
$H^{g}_{\pi}(x,\,\xi,\,t)=\frac{1}{\bar{P}^{+}}\int\frac{dz^{-}}{2\pi}e^{ix\bar{P}^{+}z^{-}}\langle\pi(P^{\prime})|\,F^{+i}\left(-\frac{z}{2}\right)F^{+i}\left(\frac{z}{2}\right)\,|\pi(P)\rangle\Big{|}_{\begin{subarray}{c}z^{+}=0\\ z^{\perp}=0\end{subarray}},$ (17)
where $F^{+\mu}(z)=\partial^{+}A^{\mu}(z)$ corresponds to the gluon field tensor in the light-cone gauge $A^{+}=0$. In the symmetric frame, the average momentum of pion $\bar{P}=\frac{1}{2}(P^{\prime}+P)$, while momentum transfer $\Delta=(P^{\prime}-P)$. The initial and final four momenta of the pion are then given by
$P$ $\equiv\left((1+\xi)\bar{P}^{+},\frac{M^{2}+\Delta_{\perp}^{2}/4}{(1+\xi)P^{+}},-\vec{\Delta}_{\perp}/2\right),$ (18)
$P^{\prime}$ $\equiv\left((1-\xi)\bar{P}^{+},\frac{M^{2}+\Delta_{\perp}^{2}/4}{(1-\xi)P^{+}},\vec{\Delta}_{\perp}/2\right),$
where $\xi=-\Delta^{+}/2\bar{P}^{+}$ and $t=\Delta^{2}$. One can derive the following relation explicitly from $\Delta^{-}$,
$-t=\frac{4\xi^{2}M^{2}+\Delta_{\perp}^{2}}{(1-\xi^{2})}\,.$ (19)
The overlap representation of the GPD in terms of the LFWFs is expressed as
$H^{g}_{\pi}(x,\,\xi,\,t)=\sum_{\lambda_{g},\lambda_{s}}\int\left[{\rm d}\mathcal{X}\,{\rm d}\mathcal{K}_{\perp}\right]\,\psi_{\lambda_{g},\lambda_{s}}^{\pi*}(x^{\prime\prime},\vec{k}_{\perp}^{\prime\prime})\psi_{\lambda_{g},\lambda_{s}}^{\pi}(x^{\prime},\vec{k}_{\perp}^{\prime}),$ (20)
where
$\left[{\rm d}\mathcal{X}\,{\rm d}\mathcal{K}_{\perp}\right]=$ $\sqrt{x^{2}-\xi^{2}}(\sqrt{1-\xi^{2}}\ )^{-1}\prod_{i=1}^{2}\left[\frac{{\rm d}x_{i}{\rm d}^{2}\vec{k}_{i\perp}}{16\pi^{3}}\right]$
×$\times\delta(x-x_{1})16\pi^{3}\delta\left(1-\sum_{i=1}^{2}x_{i}\right)\delta^{2}\left(\sum_{i=1}^{2}\vec{k}_{i\perp}\right),$ (21)
and the light-front momenta are $x^{\prime}_{1}=\frac{x_{1}+\xi}{1+\xi}$; $\vec{k}^{\prime}_{1\perp}=\vec{k}_{1\perp}+(1-x^{\prime})\frac{\vec{\Delta}_{\perp}}{2}$ for the initial gluon ($i=1$) and $x^{\prime}_{i}=\frac{x_{i}}{1+\xi};~{}\vec{k}^{\prime}_{i\perp}=\vec{k}_{i\perp}-{x_{i}^{\prime}}\frac{\vec{\Delta}_{\perp}}{2}$ for the initial spectator ($i\neq 1$), and $x^{\prime\prime}_{1}=\frac{x_{1}-\xi}{1-\xi}$; $\vec{k}^{\prime\prime}_{1\perp}=\vec{k}_{1\perp}-(1-x^{\prime\prime})\frac{\vec{\Delta}_{\perp}}{2}$ for the final gluon and $x^{\prime\prime}_{i}=\frac{x_{i}}{1-\xi};~{}\vec{k}^{\prime\prime}_{i\perp}=\vec{k}_{i\perp}+{x_{i}^{\prime\prime}}\frac{\vec{\Delta}_{\perp}}{2}$ for the final spectator. Note that, in this work, we limit ourselves to the Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) region, $\xi<x<1$, where the number of partons in the initial and the final states remains conserved.
Figure 3: The gluon GPD of the pion as a function of (a) $x$ and $-t$ [in GeV2] at fixed $\xi$, and (b) $x$ and $\xi$ at fixed $-t$ [in GeV2].
We present the pion’s gluon GPD as a function of $x$ and $-t$ at fixed $\xi=0.15$ in the left panel of Fig. 3. Physically, this corresponds to probing the pion off-forward, so the distribution is shifted toward configurations where the gluon carries slightly different longitudinal momentum after the interaction, leading to a skewed distribution. The distribution peaks when there is no transverse momentum transfer and the gluon carries less than 50% of the pion’s longitudinal momentum. As $-t$ increases, the peak shifts to higher $x$ values, and the overall magnitude decreases. At large $x$, the GPDs decay and become nearly independent of $-t$. This behavior is consistent with trends observed in gluon GPDs Lin et al. (2023); Zhang et al. (2025); Lin et al. (2025); Chakrabarti et al. (2024) and quark GPDs in both the proton Liu et al. (2024); Ji et al. (1997); Boffi et al. (2003, 2004); Mondal and Chakrabarti (2015); Mondal (2017); Freese and Cloët (2021) and the pion Kaur et al. (2018, 2020).
The right panel of Fig. 3 shows the pion’s gluon GPD as a function of $x$ and $\xi$ at fixed $t=-0.5$ GeV2. The distribution peaks at low $x$ ($<0.5$) and shifts toward higher $x$ with increasing $\xi$, while the peak height diminishes. At large $x$, the distributions again become insensitive to $\xi$. These features are model- and parton-independent and consistent with parton distributions across various QCD-inspired models Ji et al. (1997); Boffi et al. (2003, 2004); Mondal and Chakrabarti (2015); Mondal (2017); Freese and Cloët (2021); Pasquini et al. (2005); Chakrabarti and Mondal (2015).
Figure 4: The gluon PDF of the pion (a) at the model scale, and (b) at the scale $5$ GeV2 compared with the available global fits Barry et al. (2021); Novikov et al. (2020); Pasquini et al. (2023).
IV.1 Gluon PDF
[…]
Figure 5: The gluon GPD of the pion as a function of (a) $x$ and ${b_{\perp}}$ [in fm] at fixed $\xi$, and (b) $\xi$ and ${b_{\perp}}$ [in fm] at fixed $x$.
IV.2 Gluon distributions in transverse position space
We compute the gluon distribution in transverse impact-parameter space by taking the two-dimensional Fourier transform of the momentum-space GPD with respect to the transverse momentum transfer Diehl (2002); Burkardt (2003),
$H^{g}_{\pi}(x,\xi,{b}_{\perp})\;=\;\frac{1}{\left(1-\xi^{2}\right)}\int\frac{d^{2}\vec{\Delta}_{\perp}}{(2\pi)^{2}}\,e^{-i\frac{\vec{\Delta}_{\perp}}{\left(1-\xi^{2}\right)}\cdot\vec{b}_{\perp}}\,H^{g}_{\pi}\!\left(x,\xi,t=-\frac{\Delta_{\perp}^{2}+4\xi^{2}M^{2}}{1-\xi^{2}}\right),$ (22)
where the mapping $(x,\vec{\Delta}_{\perp}/(1-\xi^{2}))\to(x,\vec{b}_{\perp})$ accounts for the Lorentz-invariant skewness parameter $\xi$ and its kinematic relation to the momentum transfer $\vec{\Delta}_{\perp}$ Diehl (2003); Belitsky and Radyushkin (2005). Here, $b_{\perp}=|\vec{b}_{\perp}|$ denotes the transverse distance of the gluon relative to the center of momentum of the pion and provides a spatial image of the gluon distribution.
The left panel of Fig. 5 shows the pion’s gluon GPD in the transverse position space, when the pion transfers $15\%$ of the longitudinal momentum to its final state, i.e., at $\xi=0.15$. The right panel of Fig. 5 shows the behavior of the gluon GPD as a function of skewness $\xi$ and the transverse impact parameter ${b}_{\perp}$, for a fixed momentum fraction $x=0.6$, meaning the gluon carries 60% of the pion’s longitudinal momentum.
As $\xi$ increases, the overall magnitude of the gluon GPD decreases uniformly across all ${b}_{\perp}$, while its shape in transverse space stays essentially unchanged. This behavior reflects the fact that a larger longitudinal momentum transfer suppresses the GPD amplitude without altering its spatial extent. As $\xi=x$, the distribution vanishes, as the active gluon in the initial state carries the entire longitudinal momentum and no momentum remains to be transferred. This behavior of the pion’s gluon GPD in transverse position space is broadly consistent with trends observed in gluon GPDs Lin et al. (2023); Chakrabarti et al. (2024), as well as quark GPDs in both the proton Mondal and Chakrabarti (2015) and the pion Kaur et al. (2018).
IV.3 Gluon distributions in boost-invariant longitudinal position space
The longitudinal momentum transfer $\xi P^{+}$ is Fourier conjugate to the longitudinal distance $\tfrac{1}{2}b^{-}$, making $\xi$ conjugate to the boost-invariant longitudinal impact parameter $\sigma=\tfrac{1}{2}b^{-}P^{+}$. Fourier transforming the GPDs with respect to $\xi$ yields distributions in the boost-invariant longitudinal coordinate $\sigma$, where the DVCS amplitude exhibits a diffraction pattern Brodsky et al. (2006, 2007), reminiscent of optical diffraction. This makes the study of GPDs in longitudinal impact-parameter space particularly compelling.
The GPDs in longitudinal impact-parameter space are defined through a Fourier transform with respect to the skewness variable $\xi$:
$H(x,\sigma,t)=\int_{0}^{\xi_{\rm f}}\frac{d\xi}{2\pi}\,e^{i\xi\sigma}H(x,\xi,t),$ (23)
where the upper integration limit, denoted as $\xi_{\text{f}}$, plays a role analogous to the slit width in diffraction, setting the condition necessary for the appearance of the diffraction pattern. Since the region of interest is $\xi<x<1$, the upper limit of the integration is given by $\xi_{\rm f}=\min(x,\,\xi_{\text{max}})$. For a fixed value of $-t$, the maximum value of $\xi$ is determined by Brodsky et al. (2001):
$\xi_{\text{max}}=\frac{1}{\sqrt{1+\frac{4M^{2}}{-t}}}.$ (24)
We present the gluon GPD in longitudinal position space, $H(x,\sigma,t)$, as a function of the longitudinal impact parameter $\sigma$, at fixed $x=0.3$ and for different values of momentum transfer squared, $-t$ in the left panel of Fig. 6. The distribution exhibits a diffraction-like pattern resulting from the Fourier transform with respect to the skewness variable $\xi$. As $-t$ increases, the amplitude of the oscillations in $\sigma$ gradually decreases, indicating a suppression of the distribution. This diminishing pattern reflects reduced coherence between the initial and final states at higher momentum transfers, while the overall shape of the distribution remains largely unchanged. The observation suggests that gluonic contributions to the longitudinal structure become less prominent as the momentum transfer grows.
Figure 6: The gluon GPDs in pion in longitudinal position space with respect to $\sigma$ (a) at different values of $-t$ [in GeV2], and (b) at different values of $x$.
We also present the gluon GPDs in longitudinal position space, $H(x,\sigma,t)$, as a function of the longitudinal impact parameter $\sigma$, for a fixed value of momentum transfer squared $t=-0.5~{}\text{GeV}^{2}$ and at different values of the longitudinal momentum fraction $x$ in the right panel of Fig. 6. As $x$ increases, the diffraction pattern becomes more pronounced, with sharper peaks and a more localized distribution in $\sigma$. This indicates that gluons carrying a larger fraction of the pion’s longitudinal momentum are more tightly localized in longitudinal position space. Conversely, for smaller $x$, the distribution appears broader, reflecting a more delocalized gluonic structure. This behavior highlights the strong correlation between the momentum fraction and the spatial location of the partons encoded in the GPDs. Note that similar features have also been observed in the gluon GPDs of the proton Zhang et al. (2025), as well as in DVCS amplitudes, coordinate-space parton distributions, and Wigner distributions across a range of models Brodsky et al. (2006, 2007); Miller and Brodsky (2020); Chakrabarti et al. (2009); Manohar et al. (2011); Kumar and Dahiya (2015); Mondal and Chakrabarti (2015); Chakrabarti and Mondal (2015); Mondal (2017); Kaur et al. (2018); Maji et al. (2022); Ojha et al. (2023).
Figure 7: The GFF of pion with respect to $Q^{2}$ [in GeV2].
IV.4 Pion gravitational form factor
[…]
$A_{\pi}(Q^{2})=$ $\sum_{\lambda_{g},\lambda_{s}}\int\frac{{\rm d}x\,{\rm d}^{2}\vec{k}_{\perp}}{16\pi^{3}}\,\psi_{\lambda_{g},\lambda_{s}}\left(x,\vec{k}_{\perp}\right)\Big{[}(1-x)\,\psi_{\lambda_{g},\lambda_{s}}^{*}\left(x,\vec{k}_{\perp}-x\vec{\Delta}_{\perp}\right)$
$+x\,\psi_{\lambda_{g},\lambda_{s}}^{*}\left(x,\vec{k}_{\perp}+(1-x)\vec{\Delta}_{\perp}\right)\Big{]},$ (25)
where $Q^{2}=\Delta_{\perp}^{2}$. Figure 7 shows the GFF $A(Q^{2})$ of the pion, including separate contributions from the gluon and the spectator. At the model scale, we observe that the spectator contribution strongly dominates over the gluonic contribution. We compare our total result, $A_{\pi}(Q^{2})=A^{g}_{\pi}(Q^{2})+A^{s}_{\pi}(Q^{2})$, with the latest lattice QCD simulations Hackett et al. (2023) and find good agreement. From this result, we extract the pion’s mass (or matter) radius as $r^{2}_{\rm m}=-6A^{\prime}(0)/A(0)=(0.46~{}\mathrm{fm})^{2}$, which is very close to the corresponding lattice QCD value $r^{2}_{\rm m}=(0.41~{}\mathrm{fm})^{2}$ Hackett et al. (2023).
V Transverse momentum-dependent gluon distributions
As three-dimensional distribution functions, TMDs offer probabilistic insight into finding a parton with longitudinal momentum fraction $x$ and transverse momentum squared ${k}^{2}_{\perp}$. At equal light-front time $z^{+}=0$, the unpolarized TMD is defined through a correlator involving two gluon fields evaluated at different spacetime positions Mulders and Rodrigues (2001); Meissner et al. (2007),
$f_{1\pi}^{g}(x,{k}^{2}_{\perp})=\frac{1}{x\bar{P}^{+}}\int\frac{\mathrm{d}z^{-}\mathrm{d}^{2}\vec{z}_{\perp}}{(2\pi)^{3}}\,e^{ix\bar{P}^{+}z^{-}}\langle\pi(P)|F^{+i}\left(-\frac{z}{2}\right)F^{+i}\left(\frac{z}{2}\right)|\pi(P)\rangle\Big{|}_{z^{+}=0}.$ (26)
The gluon TMDs can be expressed in terms of overlaps of LFWFs as follows
$f_{1\pi}^{g}(x,{k}^{2}_{\perp})=\sum_{\lambda_{g},\lambda_{s}}\int\frac{{\rm d}x~{}{\rm d}\vec{k}_{\perp}}{16\pi^{3}}\Big{|}\psi^{\pi}_{\lambda_{g},\lambda_{s}}(x,\vec{k}_{\perp})\Big{|}^{2}.$ (27)
Figure 8: The $x$ multiplied TMD of gluon in pion with respect to $x$ and $k^{2}_{\perp}$ [in GeV2].
In Fig. 8, we present the gluon TMD in the pion, weighted by the momentum fraction $x$. This quantity represents the probability density of finding a gluon inside the pion with a given longitudinal momentum fraction $x$ and transverse momentum $\vec{k}_{\perp}$. The figure reveals key features of the gluonic structure of the pion in momentum space. As expected from the general behavior of gluon distributions in hadrons, the density is largest at small $x$, reflecting the dominant role of soft gluons in the nonperturbative regime of QCD. These low-$x$ gluons are abundant due to gluon radiation and the steep growth of gluon densities at small momentum fractions, a trend commonly observed in both global fits and theoretical models.
In the transverse momentum direction, the distribution exhibits a peak near $\vec{k}_{\perp}=0$, indicating that gluons are more likely to be found with low transverse momenta. The density falls off monotonically with increasing $\vec{k}_{\perp}$, eventually vanishing at large $\vec{k}_{\perp}$, as expected in a confined system where high-momentum components are suppressed. This behavior is consistent with confinement dynamics and the absence of free gluons in the hadronic spectrum.
The suppression of gluon TMDs at large $x$ and large $\vec{k}_{\perp}$ reflects the fact that high-momentum gluons are less likely to emerge from a predominantly valence-like structure at the model scale. Furthermore, the shape of the distribution can serve as an input for QCD evolution equations, such as the Collins–Soper–Sterman formalism Collins et al. (1985), to make predictions at higher scales relevant for experimental observables in gluon-sensitive processes, such as pion-induced quarkonium production, DY, and dijet production.
Figure 9: The first and second $k_{\perp}$-moment of gluon TMD in pion with respect to $x$.
Further, we compute the first and second transverse momentum moments of the pion’s gluon TMD $f_{1\pi}^{g}$, which characterize the average and the spread of the gluon’s transverse motion inside the pion. These moments are defined as Lorcé et al. (2016)
$\langle k_{\perp}^{n}\rangle^{g}_{\pi}(x)=\frac{\int\mathrm{d}^{2}\vec{k}_{\perp}\,k_{\perp}^{n}f_{1\pi}^{g}(x,k_{\perp}^{2})}{\int\mathrm{d}^{2}\vec{k}_{\perp}\,f_{1\pi}^{g}(x,k_{\perp}^{2})},$ (28)
where $n$ denotes the order of the moment. Figure 9 shows the $x$-dependence of the first ($n=1$) and second ($n=2$) transverse momentum moments. We observe that both $\langle k_{\perp}^{n}\rangle^{g}_{\pi}$ display a strong dependence on the gluon’s longitudinal momentum fraction $x$, peaking around $x=0.5$, and exhibiting approximate symmetry about this point. This behavior reflects the underlying symmetry of the pion’s internal structure in our model. The peak at intermediate $x$ suggests that gluons carrying a moderate fraction of the pion’s momentum are more likely to possess significant transverse motion. These transverse momentum moments offer valuable insight into the interplay between longitudinal and transverse dynamics of gluons and are essential inputs for TMD phenomenology and small-$x$ physics.
In addition to the $x$-dependent moments, we also compute the first and second transverse momentum moments of the gluon TMD integrated over the full range of the gluon’s longitudinal momentum fraction. These $x$-integrated moments quantify the overall transverse motion of gluons inside the pion and are defined as Lorcé et al. (2016)
$\langle k_{\perp}^{n}\rangle^{g}_{\pi}=\frac{\int_{0}^{1}\mathrm{d}x\int\mathrm{d}^{2}\vec{k}_{\perp}\,k_{\perp}^{n}f_{1\pi}^{g}(x,k_{\perp}^{2})}{\int_{0}^{1}\mathrm{d}x\int\mathrm{d}^{2}\vec{k}_{\perp}\,f_{1\pi}^{g}(x,k_{\perp}^{2})}.$ (29)
These global observables offer a complementary perspective to the $x$-dependent moments, providing an averaged measure of gluon transverse dynamics across all momentum fractions. Numerically, we find the average transverse momentum and its square to be
$\langle k_{\perp}\rangle^{g}_{\pi}=0.175~{}\text{GeV},\qquad\langle k_{\perp}^{2}\rangle^{g}_{\pi}=0.043~{}\text{GeV}^{2},$
indicating that gluons inside the pion typically carry modest transverse momentum at the model scale. These integrated moments are also useful benchmarks for comparing different theoretical models and lattice extractions of TMDs.
VI Conclusions
We have developed a light-front framework to study the gluonic structure of the pion by explicitly incorporating gluon degrees of freedom. In this model, high-energy interactions are mediated by an active gluon, while the rest of the pion is described as a spectator system with an effective mass. The mass spectrum and light-front wave functions (LFWFs) of the pion are simultaneously determined by solving the light-front holographic equation in the chiral limit for transverse dynamics and the ’t Hooft equation for longitudinal confinement. The resulting mass spectrum is consistent with experimental data, validating the predictive power of our framework.
Using the LFWFs, we computed a range of gluon distributions in the pion, including the parton distribution functions (PDFs), generalized parton distributions (GPDs) at nonzero skewness in the DGLAP region, and transverse momentum-dependent distributions (TMDs). The extracted gluon PDF shows good agreement with global fits such as JAM and xFitter, while the gravitational form factor $A(t)$ is found to be in close agreement with recent lattice QCD results. These observables, particularly the GPDs and TMDs, provide essential insights into the three-dimensional gluonic structure of the pion and serve as predictions for upcoming experiments at facilities such as the Electron-Ion Collider in USA and China, Jefferson Lab etc. These findings highlight the effectiveness of our approach in providing a coherent, relativistic, and QCD-based description of pion structure.
VII Acknowledgement
We thank Tobias Frederico for insightful and fruitful discussions. SK is supported by Research Fund for International Young Scientists, Grant No. 12250410251, from the National Natural Science Foundation of China (NSFC), and China Postdoctoral Science Foundation (CPSF), Grant No. E339951SR0. CM is supported by new faculty start up funding by the Institute of Modern Physics, Chinese Academy of Sciences, Grant No. E129952YR0. CM also thanks the Chinese Academy of Sciences Presidents International Fellowship Initiative for support via Grants No. 2021PM0023.